OpenStudy (anonymous):

Simplify. 4 square root of 2 end root plus 7 square roots of 2 end root minus 3 square roots of 2 2 square roots of 8 8 square roots of 2 8 square roots of 6 6 square roots of 8

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Factor out radical 2. \[4 \sqrt{2}+7 \sqrt{2}-3 \sqrt{2}=11 \sqrt{2} \]